OpenStudy (anonymous):

which of the following are not rational? 0.16 squared, 19 squared, 900 squared, or 100 squared?

OpenStudy (zzr0ck3r):

\[0.16^2 = (\frac{16}{100})^2\in Q\\19^2=\frac{19^2}{1}\in Q\\900^2=\frac{900^2}{1}\in Q\\100^2=\frac{100^2}{1}\in Q\]